OpenStudy (tystar123):

Select all that apply. Which steps are involved in multiplying fractions? Find a common denominator Multiply the numerators together Check to see if the product can be simplified Multiply the denominators together Convert the product to a mixed number

OpenStudy (sooobored):

\[\frac{a}{b} \times \frac{c}{d}= \frac{ a\times c}{b\times d}\]